AYYYYMANNNNN:

The table and the graph each show a different relationship between the same two variables, x and y: A table with two columns and 5 rows is shown. The column head for the left column is x, and the column head for the right column is y. The row entries in the table are 3,180 and 4,240 and 5,300 and 6,360. On the right of this table is a graph. The x axis values are from 0 to 10 in increments of 2 for each grid line. The y axis values on the graph are from 0 to 350 in increments of 70 for each grid line. A line passing through the ordered pairs 2, 70 and 4, 140 and 6, 210 and 8, 280 is drawn. How much more would the value of y be in the table than its value on the graph when x = 11? (5 points) Group of answer choices 215 110 275 150

XioGonz:

Do you have a picture?

AYYYYMANNNNN:

jhonyy9:

from this table result y = 60x so for x = 11 => y = 60*11 = ?

AYYYYMANNNNN:

im lost

jhonyy9:

calcule this value of y for x= 11 than you get it so use the rule of 3 simple and write in this way on the graph you see that to x = 2 correspond y = 70 and to x = 4 correspond y = 140 using these you can write to x = 2 correspond y = 70 to x = 11 correspond y = t t = 11*70/2 = 11*35 = ? so in this way will get the value of y when x = 11

jhonyy9:

@XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 calcule this value of y for x= 11 than you get it so use the rule of 3 simple and write in this way on the graph you see that to x = 2 correspond y = 70 and to x = 4 correspond y = 140 using these you can write to x = 2 correspond y = 70 to x = 11 correspond y = t t = 11*70/2 = 11*35 = ? so in this way will get the value of y when x = 11 \(\color{#0cbb34}{\text{End of Quote}}\) I agree! Nice (:

jhonyy9:

ty @Laylalyssa

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ty @Laylalyssa \(\color{#0cbb34}{\text{End of Quote}}\) oh ya good job :)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ty @Laylalyssa \(\color{#0cbb34}{\text{End of Quote}}\) oh ya good job :) \(\color{#0cbb34}{\text{End of Quote}}\) ty

Laylalyssa:

yw